OpenStudy (anonymous):

please help!! medal & fan z1=8(cos 2pi/3 + isin 2pi/3) z2= 0.5(cos pi/3 + isin pi/3) write the rectangular form of z1z2.

OpenStudy (anonymous):

@perl @Michele_Laino

OpenStudy (anonymous):

@paki @sammixboo

OpenStudy (anonymous):

@iambatman @EclipsedStar

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (anonymous):

@geerky42

OpenStudy (elonasushchik):

1. z1/z2 = 8 (cos π/3 + i sin π/3) = 8 (1/2 + i √3/2) = 4 + 4i √2

OpenStudy (anonymous):

is that z1? @ElonaSushchik

OpenStudy (elonasushchik):

ya

OpenStudy (anonymous):

so then how do i multiply that with z2? cause i have to multiply z1 and z2 @ElonaSushchik

OpenStudy (anonymous):

@owllover123 help me pls!!

OpenStudy (anonymous):

For two complex numbers \(\large z_1=re^{i\theta}\) and \(\large z_2=se^{i\phi}\) (where \(\large e^{i\theta}=\cos\theta+i\sin\theta\)), you have \[\large z_1z_2=rse^{i(\theta+\phi)}\] All this means is that you multiply the moduli \(r\) and \(s\), and add the arguments \(\theta\) and \(\phi\). So what's \(8\times0.5\)? What's \(\dfrac{2\pi}{3}+\dfrac{\pi}{3}\)?

OpenStudy (anonymous):

@perl @izzy529 @sammixboo